### v4

@@ I Introduction
The field of object detection has witnessed a dramatic shift in recent years, with deep learning architectures achieving unprecedented performance. However, video analysis presents a unique challenge that goes beyond the analysis of isolated frames. Unlike static images, videos possess an inherent temporal dimension, where the relationships between frames hold crucial information. Capturing these temporal dynamics is essential for a comprehensive understanding of the visual content in videos [1, 2, 3, 4, 5, 6, 7]. This is particularly important in the context of multiple object tracking (MOT), which has numerous applications in surveillance systems, intelligent robotics, and autonomous vehicles [8, 9, 10, 11].
Most recent MOT approaches leverage a tracking-by-detection paradigm [12, 13]. This involves three key steps: (1) object detection in each frame, (2) object tracking based on information from previous frames, and (3) data association using location and feature data from both detector and tracker [14]. While algorithms like SORT perform well in terms of precision and accuracy, they often suffer from ID switches and struggle with occlusions. DeepSORT addresses this with a better association metric, but real-world deployment necessitates both high accuracy and real-time performance [3].
Despite the surge in proposed MOT algorithms, achieving real-time performance remains a significant challenge. The MOTChallenge leaderboard exemplifies this inherent trade-off: faster (higher Hz) algorithms typically exhibit lower Multiple Object Tracking Accuracy (MOTA), while higher MOTA results tend to be associated with slower processing speeds.
@@ II-A Tracking-by-detection
The dominant paradigm in MOT is tracking-by-detection [15, 16, 17, 18, 19]. This approach leverages two key components: an object detector and a tracker. In each frame, the detector identifies and localizes objects. The tracker then predicts the current location of previously detected objects based on their past states. Finally, an association step links detected objects from the current frame with corresponding tracks using information from both the detector and tracker [3, 20, 21, 22, 23]. This approach offers improved accuracy compared to alternatives, as the tracker’s predictions can be continuously refined by up-to-date detection.
@@ II-B Real-time Object Tracking
Defining a universal speed threshold for real-time performance in MOT is challenging due to variations in video frame rates captured by different cameras. However, a system is generally considered real-time if its processing speed surpasses the input video frame rate [24, 25, 26]. If the processing speed falls behind, a delay accumulates, causing the system output to become out of sync with the actual video and events.
@@ II-C2 MobileNet-SSD
MobileNets are neural networks that perform efficiently. It can also be used on mobile devices and reach a fairly high accuracy [29]. SSD [30, 31, 32] uses VGG16 [33] to extract feature maps. SSD classifies and locates objects in a single forward pass. MobileNet-SSD combines SSD and MobileNets which perform high speed and relatively high accuracy. Using MobileNet SSD, the system reaches 17.5% MOTA and performs 15.604 Hz of speed.
@@ III-C Confidence Score Inference
where $i$ denotes the information at the $i$-th frame and $j$ denotes the information at $j$-th frame. The $x_{j}$ is the matrix of all the predicted bounding boxes at the $j$-th frame. The $y_{i}$ and $S_{i}$ represent the mean and covariance of detected bounding boxes. The Mahalanobis distance follows a chi-square distribution [35]. Given probability level $p$ and degrees of freedom $v$, we can calculate the distance threshold corresponding to that probability using the inverse cumulative distribution function (CDF) [37, 3, 38, 39, 40, 41]:

### v6

@@ I Introduction
The field of object detection has witnessed a dramatic shift in recent years, with deep learning architectures achieving unprecedented performance. However, video analysis presents a unique challenge that goes beyond the analysis of isolated frames. Unlike static images, videos possess an inherent temporal dimension, where the relationships between frames hold crucial information. Capturing these temporal dynamics is essential for a comprehensive understanding of the visual content in videos [1, 2, 3, 4]. This is particularly important in the context of multiple object tracking (MOT), which has numerous applications in surveillance systems, intelligent robotics, and autonomous vehicles [5, 6].
Most recent MOT approaches leverage a tracking-by-detection paradigm [7]. This involves three key steps: (1) object detection in each frame, (2) object tracking based on information from previous frames, and (3) data association using location and feature data from both detector and tracker [8]. While algorithms like SORT perform well in terms of precision and accuracy, they often suffer from ID switches and struggle with occlusions. DeepSORT addresses this with a better association metric, but real-world deployment necessitates both high accuracy and real-time performance [3].
Despite the surge in proposed MOT algorithms, achieving real-time performance remains a significant challenge [9]. The MOTChallenge leaderboard exemplifies this inherent trade-off: faster (higher Hz) algorithms typically exhibit lower Multiple Object Tracking Accuracy (MOTA), while higher MOTA results tend to be associated with slower processing speeds.
@@ II-A Tracking-by-detection
The dominant paradigm in MOT is tracking-by-detection [10]. This approach leverages two key components: an object detector and a tracker. In each frame, the detector identifies and localizes objects. The tracker then predicts the current location of previously detected objects based on their past states. Finally, an association step links detected objects from the current frame with corresponding tracks using information from both the detector and tracker [3]. This approach offers improved accuracy compared to alternatives, as the tracker’s predictions can be continuously refined by up-to-date detection.
@@ II-B Real-time Object Tracking
Defining a universal speed threshold for real-time performance in MOT is challenging due to variations in video frame rates captured by different cameras. However, a system is generally considered real-time if its processing speed surpasses the input video frame rate [11]. If the processing speed falls behind, a delay accumulates, causing the system output to become out of sync with the actual video and events.
@@ II-C2 MobileNet-SSD
MobileNets are neural networks that perform efficiently. It can also be used on mobile devices and reach a fairly high accuracy [14]. SSD [15] uses VGG16 [16] to extract feature maps. SSD classifies and locates objects in a single forward pass. MobileNet-SSD combines SSD and MobileNets which perform high speed and relatively high accuracy. Using MobileNet SSD, the system reaches 17.5% MOTA and performs 15.604 Hz of speed.
@@ III-C Confidence Score Inference
where $i$ denotes the information at the $i$-th frame and $j$ denotes the information at $j$-th frame. The $x_{j}$ is the matrix of all the predicted bounding boxes at the $j$-th frame. The $y_{i}$ and $S_{i}$ represent the mean and covariance of detected bounding boxes. The Mahalanobis distance follows a chi-square distribution [18]. Given probability level $p$ and degrees of freedom $v$, we can calculate the distance threshold corresponding to that probability using the inverse cumulative distribution function (CDF) [20, 3, 21, 22, 23, ding-24-style]: